OpenStudy (anonymous):

y varies inversely as twice x. When x = 9, y = 2. Find y when x = 3.

OpenStudy (lgbasallote):

firstly, do you know how to put "y varies inversely as twice x" in equation form?

OpenStudy (anonymous):

y = k/x I believe it is.

OpenStudy (anonymous):

Twice x bro.

OpenStudy (anonymous):

y= k/2x?

OpenStudy (anonymous):

Does that work? Do you have the answer?

OpenStudy (anonymous):

i got as far is y= k/9 I have no clue what K would stand for and Y is missing.

OpenStudy (anonymous):

It's a function. So we use the first set of points to find k, then substitute the new point into the function to get a value for y.